Can you assistance with this problem...I tried working it myself and I still did not come up with the right answer...instructions...given f(x)=-5x-3,find f(a-2)
-5(a-2)-3
-5a+10-3
-5a+3
did i do it correctly?

Answer by checkley75(3666)

f(x)=-5x-3
f(a-2)=-5(a-2)-3
f(a-2)=-5a+10-3
f(a-2)=-5a+7